OpenStudy (anonymous):

I'm having trouble...I have to factor n^2 + n + 16...

OpenStudy (anonymous):

Well that expression alone is a prime factor within the Real numbers set.

OpenStudy (anonymous):

Tho in the Complex set it has this two factors: \[n ^{2}+n+16=\left( n+0.5\left( 1+\sqrt{-47} \right) \right)\left( n+0.5\left( 1-\sqrt{-47} \right) \right)\]

OpenStudy (anonymous):

.... which are found by use of the quadratic formula

OpenStudy (anonymous):

Depending on the type or problem, you'll just want to use the "completing the square" method, and just create a perfect factor.